ISS crew work cycle shift begins with an earlier sleeptime: wake 6:32am EDT; sleep 9:32pm (Shuttle crew remaining at 10:02pm).
Crew activities aboard the ISS today centered on three major areas: (1) JLP (JEM Logistics Pressurized Module) relocation, (2) JPM (Japanese Pressurized Module) Kibo outfitting, and (3) start of JEM RMS (Robotic Manipulator System) activation & checkout.
JLP was successfully installed at its final location on the Kibo JPM at 4:04pm EDT. (After JLP/Node-2 vestibule demating and depressurization, MS1 Nyberg and FE-2-17 Chamitoff used the SSRMS (Space Station Remote Manipulator System) to grapple, unberth, transfer and reberth the JLPon Kibo's overhead port (1st stage capture 3:54pm, SSRMS wrist limped 3:58pm, 2nd stage capture with all 16 bolts 4:04pm). Karen, Greg & Aki Hoshide then latched the JPM overhead hatch via ratchet & crank handle, pressurized the connecting vestibule partially and initiated the standard vestibule gross leak check, later configuring the gear for the usual overnight fine leak check. After the installation, ISS attitude was maneuvered to the new TEA (Torque Equilibrium Attitude) which the addition of the JLP has changed. JLP was delivered on orbit by STS-123/Endeavour and docked at the Node-2 zenith port on 3/14.)
In support of the JLP transfer & berthing -
CDR Volkov connected the accelerometer in the Shuttle airlock to the IWIS RSU (Internal Wireless Integrated System/Remote Sensor Unit) and programmed/activated the system for taking structural dynamics data during the JLP relocation. (IWIS RSUs are currently deployed in the SM (Service Module), FGB, Node-1, Lab, Node-2 and Discovery. IWIS will also collect data during the upcoming DTF (Dedicated Thruster Firing) on FD10 (6/9)),
FE-2 Reisman powered up the CBCS (Centerline Berthing Camera System) which he installed yesterday at the JPM hatch, then worked with MS4 Akihiko Hoshide to configure and depressurize the JLP/Node-2 vestibule for demating,
Aki supported the demating of the Node-2 zenith CBM (Common Berthing Mechanism), petals to be closed at a later time,
CDR Mark Kelly provided camera support during the relocation,
Hoshide removed the sample port cap from the PPRV (Positive Pressure Relief Valve) in the JPM overhead hatch,
Nyberg then detached the PPRV from the JPM hatch, now no longer required, and replaced it with an MPEV (Manual Pressure Equalization Valve), and
Reisman deactivated and dismantled the no longer needed CBCS.
After yesterday's outfitting, the SSIPC (Space Station Integrated Promotion Center) activated the JPM Power Channel A last night. The activation was nominal and now both channels, A & B, are functional.
Continuing Kibo laboratory outfitting activities today -
Karen Nyberg & Greg Chamitoff rotated the ECLSS/TCS2 (Thermal Control System 2) rack down to close the manual inlet valve of the ATCS TCA (Active TCS Thermal Control Assembly) accumulator, then moved the rack back up, and
Mark Kelly & Ken Ham worked on the DMS2 (Data Management Systems 2), EPS2 (Electrical Power Systems 2) and TCS1 (Thermal Control System 1) racks in the JPM, reconfiguring each with a pivot pin and K-BAR (Knee-Brace Assembly Replacement) capture mechanism.
FE-1 Kononenko set up new Bubble dosimeters for recording radiation traces as an additional component of the RS (Russian Segment) radiation payload suite 'Matryoshka-R' (RBO-3-2), initializing and deploying the detectors. Proper function of the setup was later verified with the LULIN-5 electronics box. (A total of seven Bubble dosimeter detectors (A01-A08, A05 not used) were initialized in the Bubble dosimeter reader in the SM and positioned at their exposure locations, three at the spherical 'Phantom' unit on the DC1 panel and four in the SM starboard crew cabin on both sides of the MOSFET (metal oxide semiconductor field-effect transistor) dosimeter detector unit. The setup was photo-documented with the NIKON D2X camera and also reported to TsUP via log sheet on the BSR-TM payload channel. The complex Matryoshka payload suite is designed for sophisticated radiation studies. Note: Matryoshka is the name for the traditional Russian set of nested dolls.)
CDR Mark Kelly & PLT Ken Ham had 2.5 hrs reserved for reworking two new A/L BCMs (Airlock Battery Charge Modules) before their installation in lieu of failed units. The BCMs will be installed in slots 1 & 2 in the A/L on FD10. (The BCMs, delivered on STS-124, were discovered to have a flawed LCD (Liquid Crystal Display) circuit board with a resistor which could overheat and thereby defeat the LED (Light-Emitting Diode) backlight capability of the new units. The modification consisted in carefully opening up the BCMs (16 fasteners each), cutting the leads of the resistor and re-assembling the boxes.)
MS4 Hoshide swapped the onboard voltage/current MultiMeter with a new unit.
Garrett Reisman retrieved the RFCA FSE (Rack Flow Control Assembly/Flight Support Equipment) accumulator which on 6/4 was removed in the JPM from the LTL (Thermal Control System/Low Temperature Loop) of the TCS/EPS1 rack, installed it on the failed RFCA from Node-2 (removed 5/28) and transferred both to the Orbiter for return.
In the SM, Sergey Volkov completed the IMF (Inflight Maintenance) on the two Russian SKV air conditioners started yesterday, installing six new retainers and restoring the nominal mounting structures for the SKVs' heat exchanger-condenser fans VTK1 & VTK2. (The outfitting involved temporary deactivation of a SIGNAL-VM (DS-7A) smoke detector (#2), removing & reinstalling panels, and dismantling the old ventilator mounts for relegation to trash.)
Using the RSE1 laptop, FE-1 Kononenko spent 3.5 hrs with the new KPT-2 science payload BAR-RM, Kelvin, Ira and TTM, set up 6/4-5 for battery charging, today running tests and taking the first measurements (temperatures, relative humidity, dew point temperatures), using the RSE1 laptop. (Measurements were taken behind panels, near welds along SM structural rings and near the shell ring in the FGB for subsequent downlinking via BSR-TM channel. At the FGB pressurized shell ring, Oleg also inspected structural elements, equipment, and cable bundles for moisture, mold, or evidence of corrosion. The data, collected at locations such as feedthroughs, windows, hatches etc., are being used to get proficient in experimenting with ISS leak detection based on environmental data anomalies (temperature, humidity, and ultrasound emissions) at potential leak locations. The BAR-RM payload uses a remote infrared thermometer (Kelvin-Video), a thermohygrometer (Iva-6A), a heat-loss anemometer/thermometer (TTM-2), an ultrasound analyzer (AU-01), and a leak detector (UT2-03) to determine physical background signs of loss of ISS pressure integrity which could be indicative of leaks in the working compartments of the station. Measurements are taken in specific zones (13 in SM PkhO and 4 in DC1), both with lights, fans & ASU pump turned on and off.)
CDR Volkov conducted his second recharging of the Motorola Iridium-9505A satellite phone brought up on Soyuz 16S, a monthly routine job. (After retrieving it from its location in the TMA-12/16S descent module (BO), Sergey initiated the recharging of its lithium-ion battery, monitoring the process every 10-15 minutes as it took place. Upon completion at ~11:15am, the phone was returned inside its SSSP Iridium kit and stowed back in the BO's operational data files (ODF) container. The satphone accompanies returning ISS crews on Soyuz reentry & landing for contingency communications with SAR (Search-and-Rescue) personnel after touchdown (e.g., after an 'undershoot' ballistic reentry, as happened during the recent 15S return). The Russian-developed procedure for the monthly recharging has been approved jointly by safety officials. During the procedure, the phone is left in its fire-protective fluoroplastic bag with open flap. The Iridium 9505A satphone uses the Iridium constellation of low-Earth orbit satellites to relay the landed Soyuz capsule's GPS (Global Positioning System) coordinates to helicopter-borne recovery crews. The older Iridium-9505 phones were first put onboard Soyuz in August 2003. The newer 9505A phone, currently in use, delivers 30 hours of standby time and three hours of talk, up from 20 and two hours, respectively, on the older units.)
The CDR also conducted his second run with the Russian DZZ-2 "Diatomeya" ocean observations program, using the NIKON F-5 digital still camera with 80-200 mm lens and the HDV (high-definition) video camcorder from SM windows 7 & 8 to record color fields in oceanic waters attributed to phytoplankton blooming and observation conditions along the flight path. (June is a period of extensive phytoplankton blooming in the waters of the Northern hemisphere and is also associated with the end of a dry season and the start of a rainy season in the largest river basins of India, West Africa, and South America. Uplinked target zones were along the Honduras coastline.)
The FE-1 conducted the periodic (currently daily) checkout/verification of IP-1 airflow sensors in the various RS (Russian Segment) hatchways, including the passageways PrK (SM Transfer Compartment) -ATV, PrK -RO (SM Working Compartment), PkhO (SM Transfer Tunnel) -RO, PkhO -DC1, PkhO -FGB PGO, FGB PGO -FGB GA, FGB GA -Node-1. (This checkup is especially important when the ventilation/circulation system has to cope with a larger crew on board, currently ten persons, and one of the two Russian SKV air conditioners off (SKV-1).)
Later, Kononenko completed the periodic (about twice a month) replenishing of the Elektron oxygen generator's water supply for electrolysis, filling the KOV EDV container with water collected in CWCs (Contingency Water Containers) #1075 & #1043 from the Lab CCAA (Common Cabin Air Assembly) dehumidifier. (The 40-minute procedure is specially designed to prevent air bubbles larger than ~10 mm from getting into the BZh Liquid Unit where they could cause Elektron shutdown.)
FE-2 Reisman performed periodic 40-min. maintenance on the U.S. OGS (Oxygen Generation System), recharging the WDS (Water Delivery System) from a filled PWR (Payload Water Reservoir), leaving the PWR stowed in front of it. (Like the Russian Elektron, OGS produces O2 from water by electrolysis, dumping the also generated H2 (hydrogen) through venting.)
In the SM, Oleg completed the routine maintenance of the SOZh/ECLSS system, including ASU toilet facilities systems/replaceables. (Regular daily SOZh maintenance consists of replacement of the KTO & KBO solid waste containers, replacement of an EDV-SV waste water and EDV-U urine container, and processing U.S. condensate water as it becomes available in a filled CWC from the Lab humidifier.)
Sergey also handled the daily IMS (Inventory Management System) maintenance, updating/editing its standard 'delta file' including stowage locations, for the regular weekly automated export/import to its three databases on the ground (Houston, Moscow, Baikonur).
Shuttle crewmembers Kelly & Ham had several hours set aside for working on transfers of cargo from and to the Discovery middeck, 'choreographed' by an uplinked detailed transfer list.
With the SSRMS parked by Karen & Greg at the JEMRMS viewing position, Aki Hoshide will perform steps later tonight for relieving any mechanical strain from the Japanese RMS main arm. (This consists of turning on the arm's RTL (Robotic Laptop Terminal) and CCP (Camera Control Panel), then resetting JEU (Joint Expedited Undocking) resolvers and powering on the Ext-2 power line which provides electricity to the RMS motor mechanisms.)
Afterwards, Aki will use the laptop for checking out the arm's MUX (Data Multiplexer) and its THC (Translational Hand Controller) in the x-axis (for single-joint ops), in a partial deploy to the EVA-3 start position, and for checking out the performance of the video systems.
Earlier today, Greg Chamitoff terminated the GN2 (gaseous nitrogen) transfer from the Shuttle to the HPT (high-pressure tank) on the 'Quest' A/L by closing the manual isolation valve. Total N2 transferred: ~14 lbs.
The FE-2-17 again had about an hour for himself for general orientation (station familiarization & acclimatization) as is standard daily rule for fresh crewmembers for the first two weeks after starting station residence.
The ISS crew performed physical workout on the TVIS treadmill (CDR, FE-2) and RED resistive exercise device (FE-1).
Afterwards, Oleg copied the exercise data file to the MEC (Medical Equipment Computer) laptop for downlink, including the daily wristband HRM (Heart Rate Monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).
At ~9:51am EDT, Oleg & Sergei powered up the SM's amateur radio equipment (Kenwood VHF transceiver with manual frequency selection, headset, and power supply) and at 9:56am conducted a ham radio exchange with students at the Kursk State Polytechnic University in Kursk, central Russia.
At 12:27pm, the two Commanders, Sergey Volkov & Mark Kelly, participated in three live interactive PAO TV interviews with KMSB-TV in Tucson, AZ (Deanna Morgan), NPR (National Public Radio, Scott Simon), and APTV (Associated Press Television, Marcia Dunn).
Port SARJ Inspection Update: As a get-ahead task during yesterday's EVA-2, Fossum and Garan spent some time on the port Solar Alpha Rotary Joint, removing cover #2, inspecting the bearing surface and replacing the cover. Photos were taken and downlinked for review. The crew reported that there are no metal shavings visible in the area nor is there visible bearing damage as was discovered in the starboard SARJ bearing. No divots were seen on the bearing surface. A grease-like substance appears to be smeared on Datum A, circumferentially distributed about 1/2" from the edge, but this is not currently considered to be detrimental to the operation of the SARJ. The origin of this substance is under investigation. The port SARJ was last inspected during the Expedition 16 Stage EVA.
No CEO (Crew Earth Observation) photo targets uplinked for today.